# Robustness and classical proxy of entanglement in variants of quantum walk

### v1

@@ VII Conclusion
The simulations of the three variants (conventional, symmetric, and split-step) of QW have demonstrated that single-particle entanglement between the internal and positional degrees of freedom is robust against different types of classical randomness in both delocalized and localized regimes. The classical quantity called overlap acts as a proxy for the entanglement entropy of the walker, allowing for more experimentally accessible and achievable method of quantifying entanglement. Even though the overlap captures the behavior of entanglement in most situations, the genuine quantum nature of the entanglement is reflected by a special case with high population imbalance between the internal states which blinds the overlap. We envision advancements in characterization and manipulation of single-particle entanglement utilizing QW will find more applications in future quantum technologies.

### v2

@@ VII Conclusion
The simulations of the three variants (conventional, symmetric, and split-step) of QW have demonstrated that single-particle entanglement between the internal and positional degrees of freedom is robust against different types of classical randomness in both delocalized and localized regimes. The classical quantity called overlap acts as a proxy for the entanglement entropy of the walker, allowing for additional experimentally accessible and achievable methods of quantifying entanglement. Even though the overlap captures the behavior of entanglement in most situations, the genuine quantum nature of the entanglement is reflected by a special case with high population imbalance between the internal states which blinds the overlap. We envision advancements in characterization and manipulation of single-particle entanglement utilizing QW will find more applications in future quantum technologies.